IX is uniformly distributed over the interval [0,10], compute P{2<X<9} , P{1<X<4} and P{X<5}.


1
Expert's answer
2022-03-03T13:36:33-0500

(i)


"P(2<X<9)=\displaystyle\int_{2}^{9}\dfrac{1}{10-0}dx=[0.1x]\begin{matrix}
 9 \\
 2
\end{matrix}=0.7"

(ii)


"P(1<X<4)=\displaystyle\int_{1}^{4}\dfrac{1}{10-0}dx=[0.1x]\begin{matrix}
 4 \\
 1
\end{matrix}=0.3"

(iii)


"P(X<5)=\displaystyle\int_{0}^{5}\dfrac{1}{10-0}dx=[0.1x]\begin{matrix}
 5 \\
 0
\end{matrix}=0.5"